n=25 sample average= 96 sample standard deviation = 1.5. What is the lower bound of a 2-sided 95% confidence interval on the population standard deviation?
Answer #1

Solution :

Given that,

c = 0.95

s = 1.5

n = 25

At 95% confidence level the is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

/2,df = 0.025,24 = 39.36

and

1- /2,df = 0.975,24 = 12.40

2L = 2/2,df = 39.36

2R = 21 - /2,df = 12.40

The 95% confidence interval for is,

s (n-1) / /2,df < < s (n-1) / 1- /2,df

1.5( 25 - 1 ) / 39.36 < < 1.5( 25 - 1 ) / 12.40

1.17 < < 2.09

( 1.17 , 2.09)

The lower bound is 1.17